OpenStudy (anonymous):

Find the diffrential equation y"-8y=0

OpenStudy (turingtest):

what have you tried so far?

OpenStudy (loser66):

@TuringTest the net is not good, he may be kicked out, not intendedly.

OpenStudy (p0sitr0n):

y''=8y y'=8(1/2)y^2+C y=8(1/2)(1/3)y^3+C y=(4/3)y^3+C suppose C=0 y=0 y-(4/3)y^3=0 y(1-(4/3)y^2)=0 1=(4/3)y^2 +-sqrt(3/4)=y has 3 roots